OpenStudy (anonymous):

Select the two equations whose lines are parallel. y=6 6x+y y=-6x y-5=6(x+2)

OpenStudy (anonymous):

the second equation (6x+y) is not complete. Can you complete it? Depending on how it is completed the answer could be the second and fourth lines

OpenStudy (anonymous):

yea its 6x+y=4

OpenStudy (anonymous):

Thanks so it is second and third lines. They both have a slope of -6

OpenStudy (anonymous):

6x+y = 4 --> y = -6x + 4 y=-6x Both have same slope m = -6, (and different y intercept) Thus they are parallel